# Covering in Hamming and Grassmann Spaces: New Bounds and Reed–Solomon-Based Constructions

### v1

@@ I Introduction
Covering problems lie at the heart of several questions in coding theory, information theory, and signal processing, where one seeks structured families of objects that approximate the ambient space with minimal distortion. They are mathematically appealing in their own right, but have also found a wide range of technical applications in data compression, signal processing, clustering, sampling, and robust system design, especially in high-dimensional spaces, to name a few [torquato10, toth22]. In discrete settings, such as the Hamming space, the covering radius of a code captures its ability to represent arbitrary vectors, while in geometric settings like the Grassmann space, analogous notions arise through metrics such as the chordal distance on subspaces. Despite their shared conceptual foundation, these two domains have typically been treated with separate analytic tools and constructions. This work develops a unified perspective on covering in Hamming and Grassmann spaces, focusing on both the algorithmic task of finding nearby codewords and the quantitative task of understanding the distortion inherent to different code families.
The notion of subspace codes and their application in non-coherent communication was first developed in the seminal work of Koetter and Kschischang [KK] in the context of randomized network coding. Specifically, they developed a coding theory framework for subspace coding over finite fields. Inspired by this, a parallel framework in the field of real/complex numbers was later developed by Soleymani and Mahdavifar [Hessam22]. They introduced a new algebraic construction for one-dimensional complex subspace codes called character-polynomial (CP) codes and later extended it to higher dimensions [Hessam21].
In the Hamming space, we study the covering problem through the lens of Reed–Solomon (RS) codes [reed1960polynomial]. This well-known family of codes is not only theoretically rich but also widely used in practice, ranging from hard-disk drives to distributed storage and computing. Although various properties of RS codes have been studied extensively in the literature, they continue to inspire the development of new codes to this day. This includes the aforementioned CP codes [Hessam22], which are essentially the concatenation of a subcode of generalized Reed–Solomon (GRS) code, where a certain subset of the coefficients of the message polynomial is set to zeros, with a character function mapping the codeword coordinates to the complex unit circle. In a recent work, Gooty et al. [gooty2025precodingdesignlimitedfeedbackmiso] used CP codes to construct new precoding designs for multiple-input multiple-output (MIMO) systems with limited feedback, an important problem in wireless communications.
@@ I-A Problem Motivation
Covering can be viewed as a form of quantization in a general metric space, which is central to problems involving the approximation or representation of continuous (or large discrete) spaces using structured, finite subsets. In a canonical covering problem, one seeks to cover a metric space with the minimum number of spheres of a prescribed radius, referred to as the covering radius. This guarantees that any point in the space can be approximated by mapping it to the center of a nearby sphere with the covering radius.
A natural approach to the covering problem is to leverage well-established code families in the underlying metric space that possess strong minimum-distance properties, while shifting attention to the algorithmic challenge of quantization. Specifically, the goal is to design efficient procedures that map an arbitrary element of the space to a nearby codeword, without requiring optimal (nearest-neighbor) decoding. In the Hamming space, GRS codes are known to meet the redundancy bound [Huffman03, Corollary 11.1.3], implying that their worst-case covering radius is large and therefore unattractive from a classical covering perspective. To the best of our knowledge, this remains the only general covering result available for RS codes.
@@ I-B Related Work
The theory of covering codes has a rich history in classical information theory. Early work in this area includes characterizing upper and lower bounds on the covering radius and its computational complexity [Cohen85, McLoughlin84], bounds on the minimal size of a code of a given covering radius [Cohen86, vanLint88], the complexity of bounding the covering radius of a binary code [Cohen97], the relationship between the covering radius of a code and its subcodes [Brualdi98], the covering radius of extended Reed–Solomon codes [Dur94, Dur91], among others. We refer the reader to the extensive list of references following [Huffman03, Chapter 15]. More recently, there has been a renewed interest in determining the covering radius of various classes of codes, including Melas codes [shi2022covering], Zetterberg type codes [shi2023covering, shi2025determining], Euclidean codes [shi2025covering], and extended algebraic-geometry codes [zhu2025dual]. However, none of these recent works studied the problem of designing efficient covering algorithms and only focused on determining the covering radius. Related works in the subspace domain include asymptotic bounds and constructions [barg2002bounds, blackburn2011asymptoticbehaviorgrassmanniancodes], covering bounds in Grassmann spaces over finite fields [qian2022coveringgrassmanniancodesbounds], efficient decoding algorithms for CP codes [riasat2024decodinganalogsubspacecodes] and, more recently, sensing subspace codes [mahdavifar2024subspace, gooty2025efficientdecoderssensingsubspace].
@@ I-C Our Contribution
In this paper, we develop a unified algorithmic and information-theoretic framework for covering in both Hamming and Grassmann spaces. On the algorithmic side, we propose an efficient puncturing-based covering algorithm for GRS codes. The algorithm operates by successively puncturing the code and invoking an off-the-shelf decoder on the resulting shortened codes, until a nearby codeword is identified. We show that this procedure efficiently, i.e., with polynomial-time complexity, finds, for any input vector, a codeword within the covering radius of the input vector. We analyze the average-case performance of the proposed algorithm by characterizing the average number of punctures required for success and by estimating the resulting average covering radius, defined as the expected distance between the input vector and the selected codeword. In the Hamming space, we derive theoretical bounds on the average number of punctures and partially resolve a conjecture posed in the preliminary version of this work [riasat2025covering, Conjecture 1]. These results demonstrate that, despite their poor worst-case covering guarantees, GRS codes exhibit favorable average covering behavior.
@@
From an information-theoretic perspective, we formalize the notion of average covering radius as a natural measure of average distortion when a code is used for quantization. Building on one-shot rate-distortion results of Elkayam and Feder [elkayam20], we derive explicit non-asymptotic random-coding bounds on the average covering radius in both Hamming and Grassmann spaces. A numerical analysis of CRS codes over prime fields suggests that, at extreme high rates, these algebraic constructions asymptotically achieve average covering radii within a constant factor of the random-coding bound in the one-dimensional Grassmann space. Moreover, numerical results in the Hamming space demonstrate that RS-based codes frequently outperform random codebooks with respect to the average covering radius.
@@ II-B Chordal Distance and Subspace Code
Soleymani and Mahdavifar [Hessam22] introduced a new class of subspace codes called character-polynomial (CP) codes defined as follows. First, we recall the definition of a character.
@@
Definition 3 (CP Code [Hessam22, Definition 6]).
@@
Note that the structure of the message space ${\mathcal{F}}_{p}(k,q)$ is crucial for orthogonality and distance properties that make CP codes useful for packing subspaces in $G_{1,n}({\mathbb{C}})$ [Hessam22, Hessam21]. Recall that in the sphere packing problem, the goal is to fit in the space pairwise disjoint spheres centered at the codewords. This is fundamentally related to decoding, where, given a point in the space, the objective is to uniquely determine its nearest codeword. In particular, given a packing of the space where the spheres have radius $\tau<d/2$, where $d$ is the minimum distance of the code, one can uniquely decode any message from $\tau$ errors, and the goal is to maximize the decoding radius $\tau$. The dual problem to sphere packing is sphere covering, which is the subject of our discussion in this work. In the sphere covering problem, the goal is to cover the space using minimally overlapping spheres. This is fundamentally related to quantization, which is the process of approximating any given point in the space by its closest codeword.
@@ III-A Average Covering Radius
The (Hamming) covering radius of a block code ${\mathcal{C}}\subseteq{\mathbb{F}}_{q}^{n}$ is defined as follows [Huffman03]:
@@ III-B Average Distortion of Random Codes
In the single-shot approach to the channel coding (or lossy compression) problem, the goal is to control a random object by providing achievable bounds without any assumption on the object itself. In particular, let $\mathcal{X}$} be a set of input symbols and $\mathcal{Y}$} a set of reproduction symbols. Given a code ${\mathcal{C}}\subseteq{\mathcal{Y}}$ and distribution ${\mathcal{D}}_{{\mathcal{X}}}$ on $\mathcal{X}$}, Elkayam and Feder [elkayam20] defined the average distortion associated with $\mathcal{C}$} as follows:
@@
where $d:{\mathcal{X}}\times{\mathcal{Y}}\to{\mathbb{R}}^{+}$ is the distortion function. They gave the following explicit formula for the average distortion of a random code ${\mathcal{C}}_{{\mathcal{Y}}}:=\{Y_{j}\sim{\mathcal{D}}_{{\mathcal{Y}}}:1\leq j\leq M\}$ consisting of $M$ codewords $Y_{j}$ drawn independently from a distribution ${\mathcal{D}}_{{\mathcal{Y}}}$ on ${\mathcal{P}}({\mathcal{Y}})$ [elkayam20, Theorem 2]:
@@
  • $\operatorname{tr}(P_{x}P_{Y}){}=\operatorname{tr}(P_{Qx}P_{QY})=|\hat{Y}_{1}|^{2}\sim{\rm Beta}(1,n-1)$ by unitary invariance, for any fixed $x:=\langle\hat{x}\rangle\in{\mathcal{Y}}$, $Y$ Haar-uniform on $\mathcal{Y}$}, and unitary $Q$ with $Q\hat{x}=\hat{e}_{1}$, so that the distributions of $Y$ and $QY:=\langle\hat{Y}\rangle$ are equal [kus88].
@@
It is important to note that the conclusions of Theorem 1 and Theorem 2 do not depend on the distribution ${\mathcal{D}}_{{\mathcal{X}}}$ over which the average covering radius of the random code ${\mathcal{C}}_{{\mathcal{Y}}}$ is defined.
@@ IV Covering Using GRS Codes
Decoding algorithms have been the subject of study for many decades, and several decoding algorithms for GRS codes are known [Huffman03, Guruswami06]. In contrast, to the best of our knowledge, no explicit covering algorithms are known for GRS codes. Note that covering involves solving a more intricate minimax problem, which is computationally harder due to the need to analyze the worst-case scenario over the entire space. And, although one might use decoding algorithms for covering, their performance is bound to be subpar as spheres of radius $\tau<d/2$ do not cover the space well. Therefore, as the first algorithm of its kind to the best of our knowledge, we propose Algorithm 1 and analyze its performance in this section.
@@ IV-A Algorithm
Given an $[n,k,d]$ GRS code $\mathcal{C}$} and an input vector $y$ in its ambient space, the algorithm first tries to decode $y$. If the decoder is successful, then we have found the closest codeword to $y$. If not, it repeatedly (up to d1d-1 times) punctures the code at the last coordinate and tries to decode the corresponding substring of $y$ until the decoder is successful, when it finally returns the re-encoding of the decoder output in the original code. Note that in an actual implementation of the algorithm, it may be desirable to have the message polynomial returned directly instead of its re-encoding, as we will see in Section V-B.
@@ IV-B Improvement Using List Decoding
Consider the following modification of Algorithm 1, where $\operatorname{GRS-decode}$ is the Guruswami–Sudan list decoder (GS) [Guruswami06] and returns the best candidate in the list (i.e., the codeword with the shortest Hamming distance from $y$). There are several advantages of doing this. First, it was shown by McEliece [McEliece03, McEliece03-1] that GS almost always returns a list of size $1$, whereby implying that it is essentially a unique decoder. Next, the decoding radius of GS is given by:
@@
This implies, in particular, that Algorithm 1 will now return a $\operatorname{GRS}$ codeword $c$ within distance $\tau_{\operatorname{GS}}$ of $y$, i.e., with $d_{{\rm H}}(y,c)\leq\tau_{\operatorname{GS}}$, if such a codeword exists [McEliece03-1]. This potentially reduces the average number of punctures, thereby improving the average covering radius.
@@ IV-C Complexity Analysis
  • For unique decoding, algorithms like Berlekamp–Welch (BW) are known with $T(n)=O(n^{3})$ [ecc, Theorem 15.1.4] as well as more efficient ones with $T(n)=O(n^{2})$ and $T(n)=O(n\operatorname{polylog}(n))$ [ecc, § 15.4].
  • For list decoding, there are known implementations of $\operatorname{GS}$ with $T(n)=O(s^{4}n^{2})$ [McEliece03, § VII], [McEliece03, § VII], [Shokrollahi00], [Roth00, § V], where $s>0$ is an integer parameter known as the interpolation multiplicity.
@@ IV-D Average Number of Punctures
When running simulations, we are generally more interested in the average-case running time of an algorithm rather than the worst-case. Note that the average-case running time of Algorithm 1 depends on the average number of punctures needed for $\operatorname{GRS-decode}$ to succeed. Based on the data in Table I, we made the following conjecture in an earlier version of this work [riasat2025covering].
@@
Suppose that $\operatorname{GRS-decode}$ in Algorithm 1 is a list decoder that returns all codewords within its decoding radius. Then, uniformly averaged on all inputs $y\in{\mathbb{F}}_{q}^{n}$, the number $N_{{\rm punc}}$ of punctures needed for $\operatorname{GRS-cover}({\mathcal{C}},y)$ to succeed for an $n,k,d]_{q}$ GRS code $\mathcal{C}$} satisfies the following inequalities:
@@
Suppose that $\operatorname{GRS-decode}$ in Algorithm 1 is a unique decoder with maximum decoding radius. Then, averaged over all inputs $y\in{\mathbb{F}}_{q}^{n}$, the number $N_{{\rm punc}}$ of punctures needed for $\operatorname{GRS-cover}({\mathcal{C}},y)$ to succeed for an $n,k,d]_{q}$ GRS code $\mathcal{C}$} is given by:

### v2

@@ I Introduction
Covering problems lie at the heart of several questions in coding theory, information theory, and signal processing, where one seeks structured families of objects that approximate the ambient space with minimal distortion. They are mathematically appealing in their own right, but have also found a wide range of technical applications in data compression, signal processing, clustering, sampling, and robust system design, especially in high-dimensional spaces, to name a few [37, 38]. In discrete settings, such as the Hamming space, the covering radius of a code captures its ability to represent arbitrary vectors, while in geometric settings like the Grassmann space, analogous notions arise through metrics such as the chordal distance on subspaces. Despite their shared conceptual foundation, these two domains have typically been treated with separate analytic tools and constructions. This work develops a unified perspective on covering in Hamming and Grassmann spaces, focusing on both the algorithmic task of finding nearby codewords and the quantitative task of understanding the distortion inherent to different code families.
The notion of subspace codes and their application in non-coherent communication was first developed in the seminal work of Koetter and Kschischang [18] in the context of randomized network coding. Specifically, they developed a coding theory framework for subspace coding over finite fields. Inspired by this, a parallel framework in the field of real/complex numbers was later developed by Soleymani and Mahdavifar [36]. They introduced a new algebraic construction for one-dimensional complex subspace codes called character-polynomial (CP) codes and later extended it to higher dimensions [35].
In the Hamming space, we study the covering problem through the lens of Reed–Solomon (RS) codes [27]. This well-known family of codes is not only theoretically rich but also widely used in practice, ranging from hard-disk drives to distributed storage and computing. Although various properties of RS codes have been studied extensively in the literature, they continue to inspire the development of new codes to this day. This includes the aforementioned CP codes [36], which are essentially the concatenation of a subcode of generalized Reed–Solomon (GRS) code, where a certain subset of the coefficients of the message polynomial is set to zeros, with a character function mapping the codeword coordinates to the complex unit circle. In a recent work, Gooty et al. [14] used CP codes to construct new precoding designs for multiple-input multiple-output (MIMO) systems with limited feedback, an important problem in wireless communications.
@@ I-A Problem Motivation
Covering can be viewed as a form of quantization in a general metric space, which is central to problems involving the approximation or representation of continuous (or large discrete) spaces using structured, finite subsets. In a canonical covering problem, one seeks to cover a metric space with the minimum number of spheres of a prescribed radius, referred to as the covering radius. This guarantees that any point in the space can be approximated by mapping it to the center of a nearby sphere within the covering radius.
A natural approach to the covering problem is to leverage well-established code families in the underlying metric space that possess strong minimum-distance properties, while shifting attention to the algorithmic challenge of quantization. Specifically, the goal is to design efficient procedures that map an arbitrary element of the space to a nearby codeword, without requiring optimal (nearest-neighbor) decoding. In the Hamming space, GRS codes are known to meet the redundancy bound [17, Corollary 11.1.3], implying that their worst-case covering radius is large and therefore unattractive from a classical covering perspective. To the best of our knowledge, this remains the only general covering result available for RS codes.
@@ I-B Related Work
The theory of covering codes has a rich history in classical information theory. Early work in this area includes characterizing upper and lower bounds on the covering radius and its computational complexity [5, 24], bounds on the minimal size of a code of a given covering radius [6, 39], the complexity of bounding the covering radius of a binary code [7], the relationship between the covering radius of a code and its subcodes [4], the covering radius of extended Reed–Solomon codes [10, 9], among others. We refer the reader to the extensive list of references following [17, Chapter 15]. More recently, there has been a renewed interest in determining the covering radius of various classes of codes, including Melas codes [31], Zetterberg type codes [32, 33], Euclidean codes [34], and extended algebraic-geometry codes [40], as well as bounds on the expected Hamming distortion of linear codes over $\mathbb{F}_{2}$ and lattices in $\mathbb{R}^{n}$ [25]. However, none of these recent works studied the problem of designing efficient covering algorithms and only focused on determining the covering radius. Related works in the subspace domain include asymptotic bounds and constructions [2, 3], covering bounds in Grassmann spaces over finite fields [26], efficient decoding algorithms for CP codes [28] and, more recently, sensing subspace codes [21, 13].
@@ I-C Our Contribution
In this paper, we develop a unified algorithmic and information-theoretic framework for covering in both Hamming and Grassmann spaces. On the algorithmic side, we propose an efficient puncturing-based covering algorithm for GRS codes. The algorithm operates by successively puncturing the code and invoking an off-the-shelf decoder on the resulting shortened codes, until a nearby codeword is identified. We show that this procedure efficiently, i.e., with polynomial-time complexity, finds, for any input vector, a codeword within the covering radius of the input vector. We analyze the average-case performance of the proposed algorithm by characterizing the average number of punctures required for success and by estimating the resulting average covering radius, defined as the expected distance between the input vector and the selected codeword. In the Hamming space, we derive theoretical bounds on the average number of punctures and partially resolve a conjecture posed in the preliminary version of this work [29, Conjecture 1]. These results demonstrate that, despite their poor worst-case covering guarantees, GRS codes exhibit favorable average covering behavior.
@@
From an information-theoretic perspective, we formalize the notion of average covering radius as a natural measure of average distortion when a code is used for quantization. Building on one-shot rate-distortion results of Elkayam and Feder [11], we derive explicit non-asymptotic random-coding bounds on the average covering radius in both Hamming and Grassmann spaces. A numerical analysis of CRS codes over prime fields suggests that, at extreme high rates, these algebraic constructions asymptotically achieve average covering radii within a constant factor of the random-coding bound in the one-dimensional Grassmann space. Moreover, numerical results in the Hamming space demonstrate that RS-based codes frequently outperform random codebooks with respect to the average covering radius.
@@ II-B Chordal Distance and Subspace Code
Soleymani and Mahdavifar [36] introduced a new class of subspace codes called character-polynomial (CP) codes defined as follows. First, we recall the definition of a character.
@@
Definition 3 (CP Code [36, Definition 6]).
@@
Note that the structure of the message space ${\mathcal{F}}_{p}(k,q)$ is crucial for orthogonality and distance properties that make CP codes useful for packing subspaces in $G_{1,n}({\mathbb{C}})$ [36, 35]. Recall that in the sphere packing problem, the goal is to fit in the space pairwise disjoint spheres centered at the codewords. This is fundamentally related to decoding, where, given a point in the space, the objective is to uniquely determine its nearest codeword. In particular, given a packing of the space where the spheres have radius $\tau<d/2$, where $d$ is the minimum distance of the code, one can uniquely decode any message from $\tau$ errors, and the goal is to maximize the decoding radius $\tau$. The dual problem to sphere packing is sphere covering, which is the subject of our discussion in this work. In the sphere covering problem, the goal is to cover the space using minimally overlapping spheres. This is fundamentally related to quantization, which is the process of approximating any given point in the space by its closest codeword.
@@ III-A Average Covering Radius
The (Hamming) covering radius of a block code ${\mathcal{C}}\subseteq{\mathbb{F}}_{q}^{n}$ is defined as follows [17]:
@@ III-B Average Distortion of Random Codes
In the single-shot approach to the channel coding (or lossy compression) problem, the goal is to control a random object by providing achievable bounds without any assumption on the object itself. In particular, let $\mathcal{X}$} be a set of input symbols and $\mathcal{Y}$} a set of reproduction symbols. Given a code ${\mathcal{C}}\subseteq{\mathcal{Y}}$ and distribution ${\mathcal{D}}_{{\mathcal{X}}}$ on $\mathcal{X}$}, Elkayam and Feder [11] defined the average distortion associated with $\mathcal{C}$} as follows:
@@
where $d:{\mathcal{X}}\times{\mathcal{Y}}\to{\mathbb{R}}^{+}$ is the distortion function. They gave the following explicit formula for the average distortion of a random code ${\mathcal{C}}_{{\mathcal{Y}}}:=\{Y_{j}\sim{\mathcal{D}}_{{\mathcal{Y}}}:1\leq j\leq M\}$ consisting of $M$ codewords $Y_{j}$ drawn independently from a distribution ${\mathcal{D}}_{{\mathcal{Y}}}$ on ${\mathcal{P}}({\mathcal{Y}})$ [11, Theorem 2]:
@@
  • $\operatorname{tr}(P_{x}P_{Y}){}=\operatorname{tr}(P_{Qx}P_{QY})=|\hat{Y}_{1}|^{2}\sim{\rm Beta}(1,n-1)$ by unitary invariance, for any fixed $x:=\langle\hat{x}\rangle\in{\mathcal{Y}}$, $Y$ Haar-uniform on $\mathcal{Y}$}, and unitary $Q$ with $Q\hat{x}=\hat{e}_{1}$, so that the distributions of $Y$ and $QY:=\langle\hat{Y}\rangle$ are equal [19].
@@
It is important to note that the conclusions of Theorem 1 and Theorem 2 do not depend on the source distribution ${\mathcal{D}}_{{\mathcal{X}}}$ over which the average covering radius of the random code ${\mathcal{C}}_{{\mathcal{Y}}}$ is defined.
@@ IV Covering Using GRS Codes
Decoding algorithms have been the subject of study for many decades, and several decoding algorithms for GRS codes are known [17, 16]. In contrast, to the best of our knowledge, no explicit covering algorithms are known for GRS codes. Note that covering involves solving a more intricate minimax problem, which is computationally harder due to the need to analyze the worst-case scenario over the entire space. Although decoding algorithms may be used for covering, their performance is bound to be subpar as spheres of radius $\tau<d/2$ do not cover the space well. Therefore, as the first algorithm of its kind to the best of our knowledge, we propose Algorithm 1 and analyze its performance in this section.
@@ IV-A Algorithm
Given an $n,k,d]_{q}$ GRS code $\mathcal{C}$} and an input vector $y$ in its ambient space, the algorithm first tries to decode $y$. If the decoder is successful, then we have found the closest codeword to $y$. If not, it repeatedly (up to d1d-1 times) punctures the code at the last coordinate and tries to decode the corresponding substring of $y$ until the decoder is successful, when it finally returns the re-encoding of the decoder output in the original code. Note that in an actual implementation of the algorithm, it may be desirable to have the message polynomial returned directly instead of its re-encoding, as we will see in Section V-B.
@@ IV-B Improvement Using List Decoding
Consider the following modification of Algorithm 1, where $\operatorname{GRS-decode}$ is the Guruswami–Sudan list decoder (GS) [16] and returns the best candidate in the list (i.e., the codeword with the shortest Hamming distance from $y$). There are several advantages of doing this. First, it was shown by McEliece [22, 23] that GS almost always returns a list of size $1$, whereby implying that it is essentially a unique decoder. Next, the decoding radius of GS is given by:
@@
This implies, in particular, that Algorithm 1 will now return a $\operatorname{GRS}$ codeword $c$ within distance $\tau_{\operatorname{GS}}$ of $y$, i.e., with $d_{{\rm H}}(y,c)\leq\tau_{\operatorname{GS}}$, if such a codeword exists [23]. This potentially reduces the average number of punctures, thereby improving the average covering radius.
@@ IV-C Complexity Analysis
  • For unique decoding, algorithms like Berlekamp–Welch (BW) are known with $T(n)=O(n^{3})$ [15, Theorem 15.1.4] as well as more efficient ones with $T(n)=O(n^{2})$ and $T(n)=O(n\operatorname{polylog}(n))$ [15, § 15.4].
  • For list decoding, there are known implementations of $\operatorname{GS}$ with $T(n)=O(s^{4}n^{2})$ [22, § VII], [22, § VII], [12], [30, § V], where $s>0$ is an integer parameter known as the interpolation multiplicity.
@@ IV-D Average Number of Punctures
In practical settings, the average-case running time of an algorithm is generally of more interest rather than the worst-case. Note that the average-case running time of Algorithm 1 depends on the average number of punctures needed for $\operatorname{GRS-decode}$ to succeed. Based on the data in Table I, we made the following conjecture in an earlier version of this work [29].
@@
Suppose that $\operatorname{GRS-decode}$ in Algorithm 1 is a list decoder that returns all codewords within its decoding radius. Then, averaged uniformly on all inputs $y\in{\mathbb{F}}_{q}^{n}$, the number $N_{{\rm punc}}$ of punctures needed for $\operatorname{GRS-cover}({\mathcal{C}},y)$ to succeed for an $n,k,d]_{q}$ GRS code $\mathcal{C}$} satisfies the following inequalities:
@@
Suppose that $\operatorname{GRS-decode}$ in Algorithm 1 is a unique decoder with maximum decoding radius. Then, averaged uniformly over all inputs $y\in{\mathbb{F}}_{q}^{n}$, the number $N_{{\rm punc}}$ of punctures needed for $\operatorname{GRS-cover}({\mathcal{C}},y)$ to succeed for an $n,k,d]_{q}$ GRS code $\mathcal{C}$} is given by: